OpenStudy (anonymous):

What is the value or values that make the denominator zero? 4/x=5/2x+3

OpenStudy (anonymous):

if the denominator of that right side is (2x+3) then just solve these two equations: x = 0 and 2x+3 = 0

OpenStudy (anonymous):

first one is obvious.....

OpenStudy (anonymous):

x=0 4/(0) = infinity ??? @dpaInc with respect how can u say that it is zero???

OpenStudy (anonymous):

question asks what values make the denominator zero...

OpenStudy (anonymous):

it doesn't say to solve..

OpenStudy (anonymous):

but it is still wrong in my perspective idk ....

OpenStudy (anonymous):

ok...:)

OpenStudy (anonymous):

i'm just following instructions.. :)

OpenStudy (anonymous):

its kinda silly question

OpenStudy (anonymous):

agreed... but maybe this question will play as a restriction if she does solve the equation...

OpenStudy (anonymous):

maybe...

OpenStudy (anonymous):

@Courtney.soden , did you want this equation solved or is the question you wanted as written?

OpenStudy (anonymous):

infinity = 5/3 is the solution lol

OpenStudy (anonymous):

silly